# Evidential Transformers for Improved Image Retrieval

### v1

@@
Evidential Transformers for Improved Image Retrieval
Danilo Đorđević
ETH Zürich
Rämistrasse 101
ddordevic@ethz.ch
   Suryansh Kumar
Texas A&M University, College Station
Visual Computing, College of PVFA
suryanshkumar@tamu.edu
Sažetak
@@ 2 Methodology
$D(\mathbf{p}\mid\boldsymbol{\alpha})=\begin{cases}\frac{1}{B(\boldsymbol{\alpha})}\prod_{i=1}^{K}p_{i}^{\alpha_{i}-1}&\text{ for }\mathbf{p}\in\mathcal{S}_{K}\\ 0&\text{ otherwise }\end{cases}$ (1)
where $K$ is the number of classes in a classification task, and $\mathcal{S}_{K}$ is an $K-1$ simplex:
@@ 2.4 Distribution-Based Distances
Slika 1: Retrieval based on image embeddings. CLS token of the vision transformer (or $\alpha$ vector) is taken as the image embedding, and is used for matching between image pairs using cosine similarity.
Slika 2: Retrieval based on continuous distribution embeddings. Images are mapped to $\alpha$-parameterized Dirichlet distributions, after which matching is performed based on negative Bhattacharyya distance between them, instead of the inner product.
Slika 3: Uncertainty-based re-ranking: First retrieval is performed the regular way, by comparing the image descriptors using the cosine similarity. However, after that, a separate, evidential neural network computes uncertainties for top K results, after which re-ranking based on these uncertainties is performed.
@@ 3 Experiements and Results
Tabela 1: Comparison between the contrastively-trained DeiT-S and GC ViT-T models in category-level image retrieval tasks on the Stanford Online Products and CUB-200-2011 datasets.
@@
Slika 4: Cosine similiarities of positive and negative pairs produced by two architectures. Greater separation observed with the GC ViT indicates better processing of both local and global relations within images.
Slika 5: Number of queries exceeding a given average precision on the CUB-200-2011 dataset. GC ViT outperforms the DeiT across all average precision thresholds.
Slika 6: Number of queries exceeding a given average precision on the SOP dataset. GC ViT outperforms the DeiT across all average precision thresholds.
Tabela 2: Evidential retrieval. Recall@K on the CUB-200 2011 dataset, achieved with the GC ViT Tiny model, trained under the various evidential network frameworks. Each row corresponds to a different method of retrieval.
@@ 4 Conclusion
Literatura
  • [1] Mohamed Aly, Peter Welinder, Mario Munich, and Pietro Perona. Automatic discovery of image families: Global vs. local features. In 2009 16th IEEE International Conference on Image Processing (ICIP), pages 777–780. IEEE, 2009.
  • [2] Relja Arandjelović and Andrew Zisserman. Three things everyone should know to improve object retrieval. 2012 IEEE Conference on Computer Vision and Pattern Recognition, pages 2911–2918, 2012.
  • [3] Artem Babenko and Victor Lempitsky. Aggregating deep convolutional features for image retrieval.
  • [4] Artem Babenko, Anton Slesarev, Alexandr Chigorin, and Victor Lempitsky. Neural codes for image retrieval.
  • [5] Charles Blundell, Julien Cornebise, Koray Kavukcuoglu, and Daan Wierstra. Weight Uncertainty in Neural Networks, May 2015. arXiv:1505.05424 [cs, stat].
  • [6] Bingyi Cao, Andre Araujo, and Jack Sim. Unifying deep local and global features for image search.
  • [7] Mircea Cimpoi, Subhransu Maji, Iasonas Kokkinos, and Andrea Vedaldi. Deep filter banks for texture recognition, description, and segmentation.
  • [8] Tristan Cinquin. The bayesian transformer. Master thesis, ETH Zurich, Zurich, 2021.
  • [9] Gabriella Csurka, Christopher R. Dance, Lixin Fan, Jutta Katharina Willamowski, and Cédric Bray. Visual categorization with bags of keypoints. In European Conference on Computer Vision, 2002.
  • [10] Arthur P. Dempster. A Generalization of Bayesian Inference, pages 73–104. Springer Berlin Heidelberg, Berlin, Heidelberg, 2008.
  • [11] Alexey Dosovitskiy, Lucas Beyer, Alexander Kolesnikov, Dirk Weissenborn, Xiaohua Zhai, Thomas Unterthiner, Mostafa Dehghani, Matthias Minderer, Georg Heigold, Sylvain Gelly, Jakob Uszkoreit, and Neil Houlsby. An image is worth 16x16 words: Transformers for image recognition at scale.
  • [12] Alaaeldin El-Nouby, Natalia Neverova, Ivan Laptev, and Hervé Jégou. Training vision transformers for image retrieval.
  • [13] Yarin Gal and Zoubin Ghahramani. Bayesian convolutional neural networks with bernoulli approximate variational inference.
  • [14] Yarin Gal and Zoubin Ghahramani. Dropout as a bayesian approximation: Representing model uncertainty in deep learning. In Maria Florina Balcan and Kilian Q. Weinberger, editors, Proceedings of The 33rd International Conference on Machine Learning, volume 48 of Proceedings of Machine Learning Research, pages 1050–1059, New York, New York, USA, 20–22 Jun 2016. PMLR.
  • [15] Jakob Gawlikowski, Cedrique Rovile Njieutcheu Tassi, Mohsin Ali, Jongseok Lee, Matthias Humt, Jianxiang Feng, Anna Kruspe, Rudolph Triebel, Peter Jung, Ribana Roscher, Muhammad Shahzad, Wen Yang, Richard Bamler, and Xiao Xiang Zhu. A survey of uncertainty in deep neural networks.
  • [16] Ali Hatamizadeh, Hongxu Yin, Greg Heinrich, Jan Kautz, and Pavlo Molchanov. Global context vision transformers.
  • [17] Taejong Joo, Uijung Chung, and Min-Gwan Seo. Being bayesian about categorical probability.
  • [18] Berk Kaya, Suryansh Kumar, Carlos Oliveira, Vittorio Ferrari, and Luc Van Gool. Uncertainty-aware deep multi-view photometric stereo. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, pages 12601–12611, 2022.
  • [19] Y. Lecun, L. Bottou, Y. Bengio, and P. Haffner. Gradient-based learning applied to document recognition. Proceedings of the IEEE, 86(11):2278–2324, 1998.
  • [20] Ce Liu, Suryansh Kumar, Shuhang Gu, Radu Timofte, and Luc Van Gool. Single image depth prediction made better: A multivariate gaussian take. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, pages 17346–17356, 2023.
  • [21] Ce Liu, Suryansh Kumar, Shuhang Gu, Radu Timofte, and Luc Van Gool. Va-depthnet: A variational approach to single image depth prediction. In The Eleventh International Conference on Learning Representations, 2023.
  • [22] Ce Liu, Suryansh Kumar, Shuhang Gu, Radu Timofte, Yao Yao, and Luc Van Gool. Stereo risk: A continuous modeling approach to stereo matching. arXiv preprint arXiv:2407.03152, 2024.
  • [23] Ze Liu, Yutong Lin, Yue Cao, Han Hu, Yixuan Wei, Zheng Zhang, Stephen Lin, and Baining Guo. Swin transformer: Hierarchical vision transformer using shifted windows.
  • [24] Ziwei Liu, Ping Luo, Shi Qiu, Xiaogang Wang, and Xiaoou Tang. Deepfashion: Powering robust clothes recognition and retrieval with rich annotations. In Proceedings of IEEE Conference on Computer Vision and Pattern Recognition (CVPR), June 2016.
  • [25] David G. Lowe. Distinctive image features from scale-invariant keypoints. International Journal of Computer Vision, 60(2):91–110, 2004.
  • [26] Andrey Malinin and Mark Gales. Predictive uncertainty estimation via prior networks.
  • [27] José Mena, Oriol Pujol, and Jordi Vitrià. Dirichlet uncertainty wrappers for actionable algorithm accuracy accountability and auditability, Dec. 2019. arXiv:1912.12628 [cs, stat].
  • [28] Arsalan Mousavian and Jana Kosecka. Deep convolutional features for image based retrieval and scene categorization.
  • [29] James Philbin, Ondrej Chum, Michael Isard, Josef Sivic, and Andrew Zisserman. Object retrieval with large vocabularies and fast spatial matching. In 2007 IEEE conference on computer vision and pattern recognition, pages 1–8. IEEE, 2007.
  • [30] Filip Radenović, Giorgos Tolias, and Ondřej Chum. Fine-tuning cnn image retrieval with no human annotation. IEEE transactions on pattern analysis and machine intelligence, 41(7):1655–1668, 2018.
  • [31] Filip Radenović, Giorgos Tolias, and Ondřej Chum. CNN image retrieval learns from BoW: Unsupervised fine-tuning with hard examples.
  • [32] Filip Radenović, Giorgos Tolias, and Ondřej Chum. Fine-tuning CNN image retrieval with no human annotation.
  • [33] T. W. Rauber, A. Conci, T. Braun, and K. Berns. Bhattacharyya probabilistic distance of the Dirichlet density and its application to Split-and-Merge image segmentation. In 2008 15th International Conference on Systems, Signals and Image Processing, pages 145–148, Bratislava, Slovakia, June 2008. IEEE.
  • [34] Ali Sharif Razavian, Hossein Azizpour, Josephine Sullivan, and Stefan Carlsson. CNN features off-the-shelf: an astounding baseline for recognition.
  • [35] Karthik Abinav Sankararaman, Sinong Wang, and Han Fang. BayesFormer: Transformer with uncertainty estimation.
  • [36] Murat Sensoy, Lance Kaplan, and Melih Kandemir. Evidential deep learning to quantify classification uncertainty.
  • [37] Sivic and Zisserman. Video google: A text retrieval approach to object matching in videos. In Proceedings ninth IEEE international conference on computer vision, pages 1470–1477. IEEE, 2003.
  • [38] Hyun Oh Song, Yu Xiang, Stefanie Jegelka, and Silvio Savarese. Deep metric learning via lifted structured feature embedding. In IEEE Conference on Computer Vision and Pattern Recognition (CVPR), 2016.
  • [39] Fuwen Tan, Jiangbo Yuan, and Vicente Ordonez. Instance-level image retrieval using reranking transformers.
  • [40] Hugo Touvron, Matthieu Cord, and Hervé Jégou. Deit iii: Revenge of the vit. In Computer Vision – ECCV 2022: 17th European Conference, Tel Aviv, Israel, October 23–27, 2022, Proceedings, Part XXIV, page 516–533, Berlin, Heidelberg, 2022. Springer-Verlag.
  • [41] C. Wah, S. Branson, P. Welinder, P. Perona, and S. Belongie. The Caltech-UCSD Birds-200-2011 Dataset. Technical Report CNS-TR-2011-001, California Institute of Technology, 2011.
  • [42] Andrew Zhai and Hao-Yu Wu. Classification is a Strong Baseline for Deep Metric Learning, Aug. 2019. arXiv:1811.12649 [cs].

### v2

@@
Abstract
@@ 2 Methodology
[…]
where $K$ is the number of classes in a classification task, and $\mathcal{S}_{K}$ is an K1K-1 simplex:
@@ 2.4 Distribution-Based Distances
Figure 1: Retrieval based on image embeddings. CLS token of the vision transformer (or $\alpha$ vector) is taken as the image embedding, and is used for matching between image pairs using cosine similarity.
Figure 2: Retrieval based on continuous distribution embeddings. Images are mapped to $\alpha$-parameterized Dirichlet distributions, after which matching is performed based on negative Bhattacharyya distance between them, instead of the inner product.
Figure 3: Uncertainty-based re-ranking: First retrieval is performed the regular way, by comparing the image descriptors using the cosine similarity. However, after that, a separate, evidential neural network computes uncertainties for top K results, after which re-ranking based on these uncertainties is performed.
@@ 3 Experiements and Results
Table 1: Comparison between the contrastively-trained DeiT-S and GC ViT-T models in category-level image retrieval tasks on the Stanford Online Products and CUB-200-2011 datasets.
@@
Figure 4: Cosine similiarities of positive and negative pairs produced by two architectures. Greater separation observed with the GC ViT indicates better processing of both local and global relations within images.
Figure 5: Number of queries exceeding a given average precision on the CUB-200-2011 dataset. GC ViT outperforms the DeiT across all average precision thresholds.
Figure 6: Number of queries exceeding a given average precision on the SOP dataset. GC ViT outperforms the DeiT across all average precision thresholds.
Table 2: Evidential retrieval. Recall@K on the CUB-200 2011 dataset, achieved with the GC ViT Tiny model, trained under the various evidential network frameworks. Each row corresponds to a different method of retrieval.